# One-Prompt Strikes Back: Sparse Mixture of Experts for Prompt-based Continual Learning

### v1

@@ 4.2 Experimental Results
Performance with Different Pre-training Paradigms. In line with prior work (wang2023hierarchical, ; jiao2024vector, ), we conduct experiments on ImageNet-R using two self-supervised pre-training paradigms: iBOT-1K (zhou2021ibot, ) and DINO-1K (caron2021emerging, ). The results, presented in Table 2, demonstrate that SMoPE surpasses all other prompt-based continual learning baselines, highlighting the generalizability and robustness of SMoPE’s design across different pre-training paradigms.
@@ 5 Discussion and Conclusion
This supplementary material provides: a theoretical analysis of prompt estimation rates when using prompt-attention score aggregation (Appendix A); the detailed training algorithm for SMoPE (Appendix B); additional experimental details (Appendix C) and results (Appendix D).
@@ Appendix A Theoretical Analysis of Prompt-Attention Score Aggregation
As established in Section 3.2, the MoE models $\hat{{\bm{h}}}_{l,1},\dots,\hat{{\bm{h}}}_{l,N}$ in each attention head share a common structure of experts and score functions. Therefore, to simplify our analysis while maintaining rigor, we focus on the first head (i.e., $l=1$) and the first row of its attention matrix (i.e., $i=1$). Within this simplified setting, we present a regression-based framework to analyze the convergence of prompt expert estimation in an MoE model for prefix tuning. This framework has previously been used to study the asymptotic behavior of other MoE-based parameter-efficient fine-tuning methods, namely, low-rank adaptation [40] and LLaMA-Adapter [9]. We begin by defining the problem setting with prompt-attention score aggregation.
@@
Here, $G^{*}=\sum_{j^{\prime}=1}^{N_{p}}\exp(\beta_{0j^{\prime}}^{*})\,\delta_{(\beta_{1j^{\prime}}^{*},\eta_{j^{\prime}}^{*})}$ represents the mixing measure, i.e., a weighted sum of Dirac measures associated with the parameters $(\beta_{1j^{\prime}}^{*},\beta_{0j^{\prime}}^{*},\eta_{j^{\prime}}^{*})$ of the prompt experts. The matrix $B_{j}^{0}$ plays the same role as the matrix $\frac{E_{1}^{\top}W_{1}^{Q}{W_{1}^{K}}^{\top}E_{j}}{\sqrt{d_{v}}}$ in the score function $s_{1,j}({\bm{X}})$, while the vector $\beta_{1j^{\prime}}^{*}$ and the matrix $W$, respectively, correspond to the components $\frac{\tilde{E}^{\top}W_{1}^{Q}{W_{1}^{K}}^{\top}{\bm{p}}^{K}_{j^{\prime}}}{\sqrt{d_{v}}}$ used in the modified score function $\tilde{s}_{j^{\prime}}({\bm{X}})$. The expert functions $h({\bm{X}},\eta_{j}^{0})$ correspond to the pre-trained experts $f_{j}({\bm{X}})$, while $h({\bm{X}},\eta_{j^{\prime}}^{*})$ correspond to the new prompt experts $f_{N+j^{\prime}}({\bm{X}})$. In general, the expert functions can have flexible parametric forms, not restricted to the simple forms used in the linear-gating-prefix MoE model [23].
Least Squares Estimation. We use the least squares method [41] to estimate the unknown parameters $(\beta_{0j^{\prime}}^{*},\beta_{1j^{\prime}}^{*},\eta_{j^{\prime}}^{*})_{j^{\prime}=1}^{N_{p}}$, or, equivalently, the true mixing measure $G^{*}$. The estimator, denoted by $\widehat{G}_{n}$, is defined as:
@@
Identifiable property. We now establish the identifiability of the MoE regression function $g_{G}$. Specifically, we show that if $g_{G}({\bm{X}})=g_{G_{*}}({\bm{X}})$ for almost every $\bm{X}$}, then the measures must be identical, i.e., $G\equiv G_{*}$. For clarity, we define
@@
To bound the complexity of this function class, [41] introduced the functional
@@
where $H_{B}(\cdot)$ denotes the bracketing entropy of $\mathcal{R}_{{N^{\prime}_{p}}}(\Theta,t)$ with respect to the ${N_{p}}^{2}(\mu)$ norm, and $a\vee b$ means $\max\{a,b\}$. By adapting the proof techniques of Theorems 7.4 and 9.2 from [41], we can establish the following lemma:
@@
Let $\tau\leq\varepsilon$ and consider a $\zeta$-cover $\{\pi_{1},\dots,\pi_{\bar{N}}\}$ of $\mathcal{R}_{{N^{\prime}_{p}}}(\Theta)$ under the $N_{p}$$^{\infty}$ norm, where $\bar{N}:=N(\zeta,\mathcal{R}_{{N^{\prime}_{p}}}(\Theta),\|\cdot\|_{{L}^{\infty}})$. Construct brackets $L_{i}({\bm{X}}),U_{i}({\bm{X}})]$ for each $i\in[\bar{N}]$ by
$L_{i}({\bm{X}}):=\max\{\pi_{i}({\bm{X}})-\zeta,0\},\quad U_{i}({\bm{X}}):=\max\{\pi_{i}({\bm{X}})+\zeta,M\}.$
By construction, $\mathcal{R}_{{N^{\prime}_{p}}}(\Theta)\subseteq\bigcup_{i=1}^{\bar{N}}[L_{i},U_{i}]$ and the width satisfies $U_{i}-L_{i}\leq\min\{2\zeta,M\}$. Thus, we have
$\|U_{i}-L_{i}\|^{2}=\int(U_{i}-L_{i})^{2}\,d\mu({\bm{X}})\leq 4\zeta^{2},$
which yields that $\|U_{i}-L_{i}\|\leq 2\zeta$. Hence, by the definition of bracketing entropy,
@@
Thus, it suffices to provide an upper bound for the covering number $\bar{N}$. Let us denote $\Delta:=\{(\beta_{1},\beta_{0})\in\mathbb{R}^{Nd\times Nd}\times\mathbb{R}^{Nd}\times\mathbb{R}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$ and $\Omega:=\{\eta\in\mathbb{R}^{q}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$. Since $\Theta$ is a compact set, $\Delta$ and $\Omega$ are also compact. Therefore, there exist $\zeta$-covers $\Delta_{\zeta}$ and ${\Omega}_{\zeta}$ for $\Delta$ and $\Omega$, respectively. It can be validated that
@@
Above, $\overline{\eta}_{i}\in{\Omega}_{\zeta}$ is the closest to $\eta_{i}$ in that set and $(\overline{\beta}_{1i},\overline{\beta}_{0i})\in\Delta_{\zeta}$ is the closest to $(\beta_{1i},\beta_{0i})$ in that set. Then, it follows that
@@
The second inequality follows from the fact that softmax weights are bounded by one. The third inequality occurs since the expert $h({\bm{X}},\eta)$ is Lipschitz continuous with respect to $\eta$ with some Lipschitz constant $L_{1}({\bm{X}})>0$. Denote
$M:$ $=\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{{N^{\prime}_{p}}}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\beta_{0j^{\prime}}),$
$\overline{M}:$ $=\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{{N^{\prime}_{p}}}\exp(\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}}+\overline{\beta}_{0j^{\prime}}).$
@@
$\|g_{\check{G}}-g_{\overline{G}}\|_{{L}^{\infty}}$ $=\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}\Bigg|\frac{1}{M}\Big(\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{{N^{\prime}_{p}}}\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})h({\bm{X}},\overline{\eta}_{j})\Big)$
$\quad-\frac{1}{\overline{M}}\Big(\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{{N^{\prime}_{p}}}\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})h({\bm{X}},\overline{\eta}_{j})\Big)\Bigg|$
$\leq\Big|\frac{1}{M}-\frac{1}{\overline{M}}\Big|\cdot\sum_{i=1}^{N}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Big|\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big|$
$\quad+\sum_{j=1}^{{N^{\prime}_{p}}}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Bigg|\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{M}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{M}}\Bigg|\cdot|h({\bm{X}},\overline{\eta}_{j})|.$ (49)
@@
$\frac{1}{M}-\frac{1}{\overline{M}}\lesssim|M-\overline{M}|$ $\leq\sum_{j^{\prime}=1}^{{N^{\prime}_{p}}}\Big|\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\beta_{0j^{\prime}})-\exp(\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}}+\overline{\beta}_{0j^{\prime}})\Big|$
@@
Consequently, it follows that
$\Big|\frac{1}{M}-\frac{1}{\overline{M}}\Big|\cdot\sum_{i=1}^{N}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Big|\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big|\lesssim\zeta.$ (50)
@@
$\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{M}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{M}}$
$=$ $~\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})\Big(\frac{1}{M}-\frac{1}{\overline{M}}\Big)+\frac{1}{\overline{M}}\Big[\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})-\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})\Big].$
Recall that the parameter space is compact and the input space are bounded, then we have
$\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})\Big(\frac{1}{M}-\frac{1}{\overline{M}}\Big)\lesssim\frac{1}{M}-\frac{1}{\overline{M}}\lesssim\zeta,$
$\frac{1}{\overline{M}}\Big[\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})-\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})\lesssim(B+1)\zeta\lesssim\zeta,$
implying that
$\sum_{j=1}^{{N^{\prime}_{p}}}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Bigg|\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{M}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{M}}\Bigg|\cdot|h({\bm{X}},\overline{\eta}_{j})|\lesssim\zeta~\sum_{j=1}^{{N^{\prime}_{p}}}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~|h({\bm{X}},\overline{\eta}_{j})|\lesssim\zeta.$ (51)
It follows from Equations (49), (50) and (51) that $\|g_{\check{G}}-g_{\overline{G}}\|_{{L}^{\infty}}\lesssim\zeta$. By applying the triangle inequality, we have
@@
The definition of the covering number gives that
@@
Putting the results of Equations (48) and (52), we obtain $H_{B}\!\left(2\zeta,\mathcal{R}_{N^{\prime}_{p}}(\Theta),\|\cdot\|_{{L}^{2}(\mu)}\right)\;\lesssim\;\log(1/\tau)$. By setting $\zeta=\varepsilon/2$, we get
@@
Denote $\Psi(\delta)=\delta\cdot[\log(1/\delta)]^{1/2}$. It can be validated that $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. It follows from Equation (53) that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{R}_{{N^{\prime}_{p}}}(\Theta,\delta))$. Additionally, we set $\delta_{n}=\sqrt{\log(n)/n}$, then there exists some universal constant $c$ such that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$. Lastly, Lemma A.4 yields the conclusion of Equation (29). Hence, the proof is completed.

### v2

@@ 4.2 Experimental Results
Performance with Different Pre-training Paradigms. In line with prior work (Wang et al., 2023; Jiao et al., 2024), we conduct experiments on ImageNet-R using two self-supervised pre-training paradigms: iBOT-1K (Zhou et al., 2021) and DINO-1K (Caron et al., 2021). The results, presented in Table 2, demonstrate that SMoPE surpasses all other prompt-based continual learning baselines, highlighting the generalizability and robustness of SMoPE’s design across different pre-training paradigms.
@@ 5 Discussion and Conclusion
Reproducibility Statement
In order to facilitate the reproduction of our empirical results, we provide detailed descriptions of the experimental setup in Section 4 and Appendix C. All datasets used in this study are publicly available, enabling full replication of our experiments.
@@
This supplementary material provides: a theoretical analysis of prompt estimation rates when using prompt-attention score aggregation (Appendix A); the detailed training algorithm for SMoPE (Appendix B); additional experimental details (Appendix C) and results (Appendix D); and a discussion on the use of large language models in this work (Appendix E).
@@ Appendix A Theoretical Analysis of Prompt-Attention Score Aggregation
As established in Section 3.2, the MoE models $\hat{{\bm{h}}}_{l,1},\dots,\hat{{\bm{h}}}_{l,N}$ in each attention head share a common structure of experts and score functions. Therefore, to simplify our analysis while maintaining rigor, we focus on the first head (i.e., $l=1$) and the first row of its attention matrix (i.e., $i=1$). Within this simplified setting, we present a regression-based framework to analyze the convergence of prompt expert estimation in an MoE model for prefix tuning. This framework has previously been used to study the asymptotic behavior of other MoE-based parameter-efficient fine-tuning methods, namely, low-rank adaptation (Truong et al., 2025) and LLaMA-Adapter (Diep et al., 2025). We begin by defining the problem setting with prompt-attention score aggregation.
@@
Here, $G^{*}=\sum_{j^{\prime}=1}^{N_{p}}\exp(\beta_{0j^{\prime}}^{*})\,\delta_{(\beta_{1j^{\prime}}^{*},\eta_{j^{\prime}}^{*})}$ represents the mixing measure, i.e., a weighted sum of Dirac measures associated with the parameters $(\beta_{1j^{\prime}}^{*},\beta_{0j^{\prime}}^{*},\eta_{j^{\prime}}^{*})$ of the prompt experts. The matrix $B_{j}^{0}$ plays the same role as the matrix $\frac{E_{1}^{\top}W_{1}^{Q}{W_{1}^{K}}^{\top}E_{j}}{\sqrt{d_{v}}}$ in the score function $s_{1,j}({\bm{X}})$, while the vector $\beta_{1j^{\prime}}^{*}$ and the matrix $W$, respectively, correspond to the components $\frac{\tilde{E}^{\top}W_{1}^{Q}{W_{1}^{K}}^{\top}{\bm{p}}^{K}_{j^{\prime}}}{\sqrt{d_{v}}}$ used in the modified score function $\tilde{s}_{j^{\prime}}({\bm{X}})$. The expert functions $h({\bm{X}},\eta_{j}^{0})$ correspond to the pre-trained experts $f_{j}({\bm{X}})$, while $h({\bm{X}},\eta_{j^{\prime}}^{*})$ correspond to the new prompt experts $f_{N+j^{\prime}}({\bm{X}})$. In general, the expert functions can have flexible parametric forms, not restricted to the simple forms used in the linear-gating-prefix MoE model (Le et al., 2024).
Least Squares Estimation. We use the least squares method (van de Geer, 2000) to estimate the unknown parameters $(\beta_{0j^{\prime}}^{*},\beta_{1j^{\prime}}^{*},\eta_{j^{\prime}}^{*})_{j^{\prime}=1}^{N_{p}}$, or, equivalently, the true mixing measure $G^{*}$. The estimator, denoted by $\widehat{G}_{n}$, is defined as:
@@
Identifiable. We now establish the identifiability of the MoE regression function $g_{G}$. Specifically, we show that if $g_{G}({\bm{X}})=g_{G_{*}}({\bm{X}})$ for almost every $\bm{X}$}, then the measures must be identical, i.e., $G\equiv G_{*}$.
For clarity, define
@@
To bound the complexity of this function class, van de Geer (2000) introduced the functional
@@
where $H_{B}(\cdot)$ denotes the bracketing entropy of $\mathcal{R}_{{N^{\prime}_{p}}}(\Theta,t)$ with respect to the ${N_{p}}^{2}(\mu)$ norm, and $a\vee b$ means $\max\{a,b\}$. By adapting the proof techniques of Theorems 7.4 and 9.2 from van de Geer (2000), we can establish the following lemma:
@@
Let $\tau\leq\varepsilon$ and consider a $\zeta$-cover $\{\pi_{1},\dots,\pi_{\bar{N}}\}$ of $\mathcal{R}_{{N^{\prime}_{p}}}(\Theta)$ under the $N_{p}$$^{\infty}$ norm, where $\bar{N}:=N(\zeta,\mathcal{R}_{{N^{\prime}_{p}}}(\Theta),\|\cdot\|_{{L}^{\infty}})$. Construct brackets ${N_{p}}_{i}({\bm{X}}),U_{i}({\bm{X}})]$ for each $i\in[\bar{N}]$ by
${N_{p}}_{i}({\bm{X}}):=\max\{\pi_{i}({\bm{X}})-\zeta,0\},\quad U_{i}({\bm{X}}):=\max\{\pi_{i}({\bm{X}})+\zeta,M\}.$
By construction, $\mathcal{R}_{{N^{\prime}_{p}}}(\Theta)\subseteq\bigcup_{i=1}^{\bar{N}}[{N_{p}}_{i},U_{i}]$ and the width satisfies $U_{i}-{N_{p}}_{i}\leq\min\{2\zeta,M\}$. Thus,
$\|U_{i}-{N_{p}}_{i}\|^{2}=\int(U_{i}-{N_{p}}_{i})^{2}\,d\mu({\bm{X}})\leq 4\zeta^{2},$
so that $\|U_{i}-{N_{p}}_{i}\|\leq 2\zeta$. Hence, by the definition of bracketing entropy,
@@
Therefore, we need to provide an upper bound for the covering number $\bar{N}$. In particular, we denote $\Delta:=\{(\beta_{1},\beta_{0})\in\mathbb{R}^{Nd\times Nd}\times\mathbb{R}^{Nd}\times\mathbb{R}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$ and $\Omega:=\{\eta\in\mathbb{R}^{q}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$. Since $\Theta$ is a compact set, $\Delta$ and $\Omega$ are also compact. Therefore, we can find $\zeta$-covers $\Delta_{\zeta}$ and ${\Omega}_{\zeta}$ for $\Delta$ and $\Omega$, respectively. We can check that
@@
Here, $\overline{\eta}_{i}\in{\Omega}_{\zeta}$ such that $\overline{\eta}_{i}$ is the closest to $\eta_{i}$ in that set, while $(\overline{\beta}_{1i},\overline{\beta}_{0i})\in\Delta_{\zeta}$ is the closest to $(\beta_{1i},\beta_{0i})$ in that set. From the above formulations, we get that
@@
Here, the second inequality occurs as the softmax weight is bounded by one, and the third inequality follows from the fact that the expert $h({\bm{X}},\cdot)$ is a Lipschitz function with some Lipschitz constant $L_{1}({\bm{X}})>0$. Next, let us denote
$D:$ $=\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{{N^{\prime}_{p}}}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\beta_{0j^{\prime}}),$
$\overline{D}:$ $=\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{{N^{\prime}_{p}}}\exp(\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}}+\overline{\beta}_{0j^{\prime}}).$
@@
$\|g_{\check{G}}-g_{\overline{G}}\|_{{L}^{\infty}}$ $=\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}\Bigg|\frac{1}{D}\Big(\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{{N^{\prime}_{p}}}\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})h({\bm{X}},\overline{\eta}_{j})\Big)$
$\quad-\frac{1}{\overline{D}}\Big(\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{{N^{\prime}_{p}}}\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})h({\bm{X}},\overline{\eta}_{j})\Big)\Bigg|$
$\leq\Big|\frac{1}{D}-\frac{1}{\overline{D}}\Big|\cdot\sum_{i=1}^{N}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Big|\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big|$
$\quad+\sum_{j=1}^{{N^{\prime}_{p}}}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Bigg|\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{D}}\Bigg|\cdot|h({\bm{X}},\overline{\eta}_{j})|.$ (49)
@@
$\frac{1}{D}-\frac{1}{\overline{D}}\lesssim|D-\overline{D}|$ $\leq\sum_{j^{\prime}=1}^{{N^{\prime}_{p}}}\Big|\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\beta_{0j^{\prime}})-\exp(\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}}+\overline{\beta}_{0j^{\prime}})\Big|$
@@
As a result, we deduce that
$\Big|\frac{1}{D}-\frac{1}{\overline{D}}\Big|\cdot\sum_{i=1}^{N}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Big|\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big|\lesssim\zeta.$ (50)
@@
$\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{D}}$
$=$ $~\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})\Big(\frac{1}{D}-\frac{1}{\overline{D}}\Big)+\frac{1}{\overline{D}}\Big[\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})-\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})\Big].$
Since both the input space and the parameter space are bounded, we have
$\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})\Big(\frac{1}{D}-\frac{1}{\overline{D}}\Big)\lesssim\frac{1}{D}-\frac{1}{\overline{D}}\lesssim\zeta,$
$\frac{1}{\overline{D}}\Big[\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})-\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})\lesssim(B+1)\zeta\lesssim\zeta,$
which yields that
$\sum_{j=1}^{{N^{\prime}_{p}}}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~\Bigg|\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{D}}\Bigg|\cdot|h({\bm{X}},\overline{\eta}_{j})|$
$\hskip 227.62204pt\lesssim\zeta~\sum_{j=1}^{{N^{\prime}_{p}}}\sup_{{\bm{X}}\in\mathcal{{\bm{X}}}}~|h({\bm{X}},\overline{\eta}_{j})|\lesssim\zeta.$ (51)
From Equations (49), (50) and (A.1), we obtain that $\|g_{\check{G}}-g_{\overline{G}}\|_{{L}^{\infty}}\lesssim\zeta$. According to the triangle inequality, we have
@@
By definition of the covering number, we deduce that
@@
Combined with Equations (48) and (52), we achieve the following result:
$H_{B}\!\left(2\zeta,\mathcal{R}_{N^{\prime}_{p}}(\Theta),\|\cdot\|_{{L}^{2}(\mu)}\right)\;\lesssim\;\log(1/\tau).$
Let $\zeta=\varepsilon/2$, then we obtain that
@@
Let $\Psi(\delta)=\delta\cdot[\log(1/\delta)]^{1/2}$, then $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Furthermore, Equation (53) indicates that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{R}_{{N^{\prime}_{p}}}(\Theta,\delta))$. In addition, let $\delta_{n}=\sqrt{\log(n)/n}$, then we get that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$ for some universal constant $c$. Finally, by applying Lemma A.4, we achieve the desired conclusion of the theorem.